OpenStudy (anonymous):

Can you show me the steps of this identity cos(3X) ?

OpenStudy (shamim):

Cos3x=4cos^3x-3cosx?

OpenStudy (shamim):

M i right?

OpenStudy (anonymous):

\[\cos(3x)=\cos(2x+x)=\cos(2x)\cos(x)-\sin(2x)\sin(x)\]\[\cos(2x)=\cos(x+x)=\cos(x)\cos(x)-\sin(x)\sin(x)=\cos^2x-\sin^2x\]\[\sin(2x)=\sin(x+x)=\sin(x)\cos(x)+\sin(x)\cos(x)=2\sin(x)\cos(x)\] So, you get,

OpenStudy (anonymous):

\[\cos(3x)=\cos(2x)\cos(x)-\sin(2x)\sin(x)\]\[\cos(3x)=(\cos^2x-\sin^2x)\cos(x)-2\sin(x)\cos(x)\sin(x)\]\[\cos(3x)=\cos^3(x)-\sin^2(x) \cos(x)-2\sin^2(x)\cos(x)\]

OpenStudy (anonymous):

THen I suppose you can play with the answer if you want to....

OpenStudy (anonymous):

yes, i've just figure the answer! Thank you guys.